OpenStudy (anonymous):

write the expressions as a complex number in standard form. -2(3+2i)

OpenStudy (mathstudent55):

Just distribute the -2 outsisde the parentheses.

OpenStudy (anonymous):

-6-4i

OpenStudy (mathstudent55):

-2(3+2i) = -2 * 3 + (-2) * 2i = -6 - 4i

OpenStudy (anonymous):

(4+3i)(1-i)

OpenStudy (mathstudent55):

Here you use FOIL.

OpenStudy (mathstudent55):

(4+3i)(1-i) = 4 * 1 + 4 * (-i) + 3i * 1 + 3i * (-i)

OpenStudy (mathstudent55):

Now do all the multiplications and simplify.

OpenStudy (anonymous):

is it 4+10i

OpenStudy (mathstudent55):

(4+3i)(1-i) = 4 * 1 + 4 * (-i) + 3i * 1 + 3i * (-i) = 4 - 4i + 3i - 3i^2 = 4 -i - (3)(-1) = 4 - i + 3 = 7 - i

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

so if you have a (1-i)(1+i) do they cancel oout

OpenStudy (mathstudent55):

You have to either multiply it out using FOIL, or remember the product of a sum and a difference where the middle terms cancel out. Let's use FOIL: (1-i)(1+i) = 1 * 1 + 1 * i - i * 1 - i * i = 1 + i - i - i^2 = 1 - (-1) = 2

OpenStudy (mathstudent55):

If you notice that (1-i)(1+i) is the product of a sum and a difference, and you remember that the product of a sum and a difference gives the difference of two squares, (a + b)(a - b) = a^2 - b^2, then you can use that: (1-i)(1+i) = 1^2 - i^2 = 1 - (-1) = 1 + 1 = 2

OpenStudy (anonymous):

i am going to show you a problem but dont help me with i just want you to tell me if its right